OpenStudy (anonymous):

The rate constant at 25C is 0.45 M/s and 0.35 M/s at 1.0C. What is the activation energy for this reaction?

OpenStudy (anonymous):

I know the formula to find this is \[\ln\left({k_2\over k_1}\right)={Ea\over R}\left({{1\over T_1}-{1\over T_2}}\right)\] but I'm not sure what R means...?

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@Mertsj @myininaya

OpenStudy (mertsj):

R is the ideal gas constant.

OpenStudy (anonymous):

and what is that?

OpenStudy (anonymous):

@AccessDenied can you help?

OpenStudy (anonymous):

@Compassionate any idea?

OpenStudy (anonymous):

wait i got it..